# appAppendix References

### v6

@@ 4.3 Does hidden dissent move the market?
In the preceding subsections, we explored various methods to measure hidden dissent within FOMC meetings, using sources such as meeting transcripts and public speeches by members. These analyses offered insights into the internal divergences of opinion among FOMC members and their relationships with economic indicators. Now, we shift our focus to understanding whether and how this hidden dissent influences financial markets.
@@
To accomplish this, we employ a MHSA-based deep learning model, a modified version of the model used earlier in this paper, to predict the hidden dissent from minutes. The model architecture, illustrated in Figure 7, is designed to effectively handle the complex language in FOMC minutes, and accurately extract key features to predict the hidden dissent level revealed in corresponding transcript. The training set is constructed by extracting content related to committee members from the minutes and labeling it with $HD_{i}$, the hidden dissent level detected from transcript $i$. Following the training procedures in Section 3, the optimal model configuration includes six MHSA modules (each with four heads), a dropout rate of 0.46, and an initial learning rate of $4.57\times 10^{-5}$. This setup achieves an average MAE of $5.835\times 10^{-2}$, an $R^{2}$ of $0.704$ on the training set, with $5.838\times 10^{-2}$ and $0.697$ on the test set, respectively, based on five-fold cross-validation. We then apply the trained model to post-2018 minutes to extend the date to the present. As shown in Figure 8 the hidden dissent level in transcripts and minutes align closely, with a correlation of $0.848$. Meanwhile, $HD_{i}^{min}$ shows a clear increase in hidden dissent during the second half of 2022, as the FOMC implemented four consecutive 75-basis-point rate hikes followed by a 50-basis-point hike, pushing its benchmark interest rate to the highest level in 15 years.161616 CNBC: Fed raises interest rates half a point to highest level in 15 years
Most existing studies predominantly focus on the sentiment conveyed through various FOMC communication channels. In this subsection, we propose that hidden dissent offers an additional layer of information beyond the general meeting sentiment and demonstrate its significant influence on financial markets. Specifically, we estimate the following model following Swanson (2021) and Gorodnichenko et al. (2023):
$Outcome_{t,t+h}=b_{0}^{(h)}+b_{1}^{(h)}HD^{min}_{t}+b_{2}^{(h)}S^{min}_{t}+error^{(h)}_{t}$
@@
Following Gorodnichenko et al. (2023), we define the outcome variables as the daily financial indicators (FI) over different horizon, $h$. We select the horizon $h=[0,15]$ where $h=0$ represents the minute release date and estimate coefficients for each horizon separately to illustrate the dynamics of the response of these financial indicators. The list of indicators can be found in Table 10 and the corresponding data is sourced from Yahoo Finance. Returns on different FIs are measured as the log of the close price at date $t+h$ minus the log of the open price at date $t$: $Outcome_{t,t+h}^{FI}=\log(FI_{t+h}^{close})-\log(FI_{t}^{open})$. Given the data availability of different financial indicators, our analysis covers the recent FOMC meetings held between 2011 and 2024. Additionally, we employ the bias-corrected and accelerated (BCa) bootstrap method to correct potential biases and construct accurate confidence intervals. [^17] We report results using the residual-based bootstrap method rather than the data-based bootstrap method. Swanson (2021) provides a detailed explanation of why the residual-based bootstrap method is preferred in the current setting. We also conduct the data-based BCa bootstrap, and the results are similar.
@@
We begin by examining the SPY index to assess the stock market’s response to hidden dissent within FOMC meetings. Our findings suggest that higher levels of hidden dissent generally lead to a decline in share prices (see Figure 9). Next, we employ the VIX index to evaluate how stock market volatility expectations react to hidden dissent. Plots in Figure 10 shows that on the minute release date, higher hidden dissent detected in the minutes significantly increases market volatility in the ensuing few days. This observation suggests that the stock market responds to newly disclosed information about hidden dissent, which is not conveyed through other communication channels on the FOMC meeting day. In contrast, sentiment information embedded in the minutes does not significantly affect stock market.
@@
We first examine how U.S. Treasury bond reacts to hidden dissent in FOMC meetings. As shown in Figure 11, hidden dissent does not substantially affect GOVT bond prices, though there is a slight upward trend indicating that higher dissent levels may put upward pressure on bond prices. However, dovish sentiment appears to drive down Treasury bond price. We then analyze the LQD index to assess the impact of hidden dissent on investment-grade corporate bonds (Figure 12). The results indicate that the effects are not statistically significant.
(3) Inflation Expectation and Interest Rate Risk
To measure changes in inflation expectations, we use the spread: $\log(\frac{P^{GOVT}{t+h,close}}{P^{GOVT}{t,open}})-\log(\frac{P^{TIP}{t+h,close}}{P^{TIP}{t,open}})$, where an increase in this spread signals a decline in expected inflation. Figure 13 shows that a higher hidden dissent level tends to increase inflation expectations, but the effect is not substantial. Additionally, dovish sentiment appears to temporarily elevate inflation expectations.
Following Gorodnichenko et al. (2023), we measure interest rate risk using the spread: $\log(\frac{P^{LQD}_{t+h,close}}{P^{LQD}_{t,open}})-\log(\frac{P^{LQDH}_{t+h,close}}{P^{LQDH}_{t,open}})$, where an increase in this spread indicates higher perceived interest rate risk. Figure 14 suggests that higher hidden dissent in FOMC meetings is associated with an increase in perceived interest rate risk. This suggests that greater hidden dissent may increase uncertainty about future monetary policy, economic growth, or market conditions, ultimately driving up interest rate risk.
(4) Foreign Exchange Market
Finally, we investigate the impact of hidden dissent on the foreign exchange market by analyzing the USD/JPY and USD/EUR exchange rates. The results indicate that hidden dissent has no significant impact on exchange rates. Meanwhile, dovish sentiment tends to drive the dollar to appreciate against the euro (see Figure 16). However, this effect is not statistically significant for USD/JPY (see Figure 15).
The above analysis demonstrates that hidden dissent, as measured through the minutes, provides additional information not captured by other FOMC communication channels and has a measurable impact on various financial markets. This influence is particularly evident in the stock market and in shaping the public’s expectations regarding future interest rate risk.
@@ 5 Conclusion
This paper has three main findings. First, hidden dissent is strongly correlated with macroeconomic variables and the education background of members, while other personal characteristics have little influence. Second, hidden dissent within FOMC meeting appears to stem from heterogeneity in both members’ preferences and their projections. Third, financial markets respond to hidden dissent, treating it as a complementary channel alongside other FOMC communication channels. Higher levels of hidden dissent tend to increase stock market volatility expectation, decrease share prices and also drives up interest rate risk.
@@ Appendix A A Brief Introduction of Deep Learning in NLP
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Results for Robustness Checks
Table B.3: Level of Hidden Dissent and FOMC Economic Projections
$HD_{i}$
$(1)$ $(2)$ $(3)$ $(4)$
$\text{Policy}_{PC1}$ $-0.1299^{**}$ $-0.1335^{**}$
$(0.0632)$ $(0.0621)$
$\text{Policy}_{PC2}$ $0.0831$
$(0.0691)$
$\text{Economy}_{PC1}$ $-0.0426^{*}$ $-0.0606^{**}$
$(0.0248)$ $(0.0308)$
$\text{Economy}_{PC2}$ $0.0066$ $0.0204$
$(0.0672)$ $(0.0655)$
$\text{Economy}_{PC3}$ $-0.0526$ $0.0270$
$(0.0723)$ $(0.0730)$
Pseudo $R^{2}$ 0.1353 0.1755 0.0839 0.1259
Log likelihood $29.9394$ $30.6291$ $41.7010$ $29.7825$
$N$ $29$ $29$ $40$ $29$
  • Note: Similar to Table 8. In this table, the disagreement level in each projection category are calculated using the absolute sum of deviations from the mean projection, instead of the median projection. ${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
@@ C.1 Hidden dissent inside FOMC members’ speeches
Utilizing the score $hd_{ij}^{speech}$ from members’ speeches, we can also construct the variable $HD_{i}^{speech}$ to measure the hidden dissent level prior to each FOMC meeting. These speeches are linked to meeting $i$ if delivered between meetings $i-1$ and $i$, and $HD_{i}^{speech}$ is calculated using the same method as $HD_{i}$. Summary statistics for these hidden dissent measures are reported in Table C.1.
@@ C.2 Do public speeches reveal hidden dissent in advance?
As discussed in subsection C.1, we feed FOMC members’ speeches into the deep learning model and generate two hidden dissent scores for each speech, $hd_{ij}^{speech,in}$ and $hd_{ij}^{speech,pre}$, depending on how to chose the reference point for chair. The first measure, $hd_{ij}^{speech,in}$, uses the chair’s words from the following meeting as the reference point. Since transcripts are subject to a 5-year delay and are not available in real time, the second measure, $hd_{ij}^{speech,pre}$, uses the chair’s speeches prior to the meeting, which are available in real time, as the reference point instead. For both measures, we compute $HD_{i}^{speech}$ for each meeting $i$ by averaging the hidden dissent scores for speeches delivered between meetings $i-1$ and $i$.

### v7

@@ 4.3 Does hidden dissent move the market?
In the preceding subsections, we explored various methods to offer insights into the internal divergences of opinion among FOMC members and their relationships with economic indicators. Now, we shift our focus to understanding whether and how this hidden dissent influences financial markets.
@@
To accomplish this, we employ a MHSA-based deep learning model, a modified version of the model used earlier in this paper, to predict the hidden dissent from minutes. The model architecture, illustrated in Figure 7, is designed to effectively handle the complex language in FOMC minutes, and accurately extract key features to predict the hidden dissent level revealed in corresponding transcript. The training set is constructed by extracting content related to committee members from the minutes and labeling it with $HD_{i}$, the hidden dissent level detected from transcript $i$. Following the training procedures in Section 3, the optimal model configuration includes six MHSA modules (each with four heads), a dropout rate of 0.46, and an initial learning rate of $4.57\times 10^{-5}$. This setup achieves an average MAE of $5.835\times 10^{-2}$, an $R^{2}$ of $0.704$ on the training set, with $5.838\times 10^{-2}$ and $0.697$ on the test set, respectively, based on five-fold cross-validation. We then apply the trained model to post-2018 minutes to extend the date to the present. As shown in Figure 8 the hidden dissent level in transcripts and minutes align closely, with a correlation of $0.848$. Meanwhile, $HD_{i}^{min}$ shows a clear increase in hidden dissent during the second half of 2022, as the FOMC implemented four consecutive 75-basis-point rate hikes followed by a 50-basis-point hike, pushing its benchmark interest rate to the highest level in 15 years. [^16] CNBC: Fed raises interest rates half a point to highest level in 15 years
While the overall sentiment (hawkish/dovish tone) of FOMC communications is known to influence markets, we propose in this section that the hidden dissent provides a separate, crucial information channel reflecting policy uncertainty. [^17] A potential concern with using generated regressors in downstream tasks, as highlighted by Battaglia et al. (2024), is that measurement error can bias coefficient estimates. The authors show this bias is governed by the parameter $\kappa=\sqrt{n}\times\mathbb{E}[1/C_{i}]$, where a small $\kappa$ indicates that the bias is likely to be minimal. In our application, the average text length ($C_{i}$) is large, leading to very small $\kappa$ values: approximately 0.011 for our transcript-based model and 0.006 for our minutes-based model. These values are substantially smaller than the 0.44 benchmark in the CEO application re-analyzed by Battaglia et al. (2024), for which they find the bias to be negligible. We therefore conclude that this issue is unlikely to be a significant concern for our estimates. We thank our anonymous referee for pointing this out. To isolate its impact, we explicitly control for sentiment in our analysis, estimate the model following Swanson (2021) and Gorodnichenko et al. (2023).
$Outcome_{t,t+h}=b_{0}^{(h)}+b_{1}^{(h)}HD^{min}_{t}+b_{2}^{(h)}S^{min}_{t}+error^{(h)}_{t}$ (3)
@@
Following Gorodnichenko et al. (2023), we define the outcome variables as the financial indicators (FI) change over different horizon, $h$. We select the horizon $h=[0,15]$ where $h=0$ represents the minute release date and estimate coefficients for each horizon separately to illustrate the dynamics of the response of these financial indicators. The list of indicators can be found in Table 11. Additionally, we employ the bias-corrected and accelerated (BCa) bootstrap method to correct potential biases and construct accurate confidence intervals. [^18] We report results using the residual-based bootstrap method rather than the data-based bootstrap method. Swanson (2021) provides a detailed explanation of why the residual-based bootstrap method is preferred in the current setting. We also conduct the data-based BCa bootstrap, and the results are similar.
@@
We begin by examining the stock market’s response to hidden dissent using daily total returns of the S&P 500 ETF (SPY). Returns are measured as the log of the close price at date $t+h$ minus the log of the open price at date $t$: $Outcome_{t,t+h}^{SPY}=\log(SPY_{t+h}^{close})-\log(SPY_{t}^{open})$. Figure 9 indicates that higher levels of hidden dissent predict a statistically significant decline in the SPY total return.
Furthermore, we employ the VIX index to evaluate how stock market volatility expectations react to hidden dissent. Plots in Figure 10 shows that on the minute release date, higher hidden dissent detected in the minutes significantly increases market volatility in the ensuing few days. This observation suggests that the stock market responds to newly disclosed information about hidden dissent, which is not conveyed through other communication channels on the FOMC meeting day. In contrast, sentiment information embedded in the minutes does not significantly affect stock market. Meanwhile, as shown in subsection 4.1, VIX has no explanatory power over hidden dissent, our current findings clearly suggest a one-way influence, reinforcing the view that markets react to FOMC communications, not vice versa.
@@
Next, we examine the reaction of U.S. Treasury yields to hidden dissent revealed in FOMC minutes. We analyze the cumulative change in yields over the 14 days following the minutes’ release, using the 10-year bond yield (DGS10Y). Overall, the results presented in Figure 11 indicate that higher hidden dissent drives up long-term bond yield changes. Meanwhile, we find that the impact of sentiment on Treasury bond is insignificant, suggesting that hidden dissent and sentiment transmit information through distinct channels.
Following Gorodnichenko et al. (2023), we further measure interest rate risk using the spread: $\log(\frac{P^{LQD}_{t+h,close}}{P^{LQD}_{t,open}})-\log(\frac{P^{LQDH}_{t+h,close}}{P^{LQDH}_{t,open}})$, where an increase in this spread indicates higher perceived interest rate risk. Figure 12 suggests that higher hidden dissent in FOMC meetings is associated with an increase in perceived interest rate risk. This suggests that greater hidden dissent may increase uncertainty about future monetary policy, economic growth, or market conditions, ultimately driving up interest rate risk.
The above analysis demonstrates that hidden dissent, as measured through the minutes, provides additional information not captured by other FOMC communication channels and has a measurable impact on various financial markets.
@@ 5 Conclusion
This paper has three main findings. First, hidden dissent is strongly correlated with macroeconomic variables and the education background of members, while other personal characteristics have little influence. Second, hidden dissent within FOMC meeting appears to stem from heterogeneity in both members’ preferences and their projections. Third, and critically, we establish hidden dissent as a distinct information channel for financial markets, separate from overall policy sentiment. By analyzing market reactions to timely minutes releases while controlling for sentiment, we show that higher hidden dissent significantly impacts both major financial markets: in the stock market, it tends to increase volatility expectations and decrease share prices; in the bond market, it drives up yields and perceived interest rate risk. This confirms that how the committee reaches a decision, specifically the degree of underlying consensus or dissent, is valuable information for market participants beyond the stated policy or general tone.
@@ Appendix A A Brief Introduction of Deep Learning in NLP
In earlier natural language processing (NLP) models, researchers often rely on manually crafted features to project text data into low-dimensional feature vectors. One representative example is the Bag of Words method. Researchers need to have sufficient domain knowledge to pick out the most relevant features, which can be time-consuming and difficult to generalize. And these NLP models typically neglect intricate interconnections between different words and sentences within an article.. Gentzkow et al. (2019) note that, the emergence of new machine learning models for NLP in computer science offers novel solutions for addressing certain economic challenges. In recent years, Deep Learning models have advanced rapidly and enabled us to surmount the constraints of previous-generation NLP models. Moreover, the introduction of Transfer Learning allows us to pre-train a complex model on extensive datasets, so we can achieve high prediction accuracy even if the target dataset is small.
The root of Deep Learning can be traced back to 1943 when Walter Pitts and Warren McCulloch created the first neural network model based on the neural networks of the human brain. Until recently, the commonly used Deep Learning models in text classification can be broadly categorized into two groups: Convolutional Neural Network (CNN) and Recurrent Neural Network (RNN) models. [^19] There are other equally important deep learning models like MLP, Autoencoder, GAN, etc. However, to our best knowledge, they are less solely used in text classification tasks. In text classification tasks, CNN models excel in detecting position-related patterns within text, while RNN models, particularly the Long Short-Term Memory (LSTM) model, exhibit robust performance in training with lengthy text. By introducing the Gated Recurrent Unit (GRU), LSTM can effectively address the vanishing gradients problem in long text classification tasks(Minaee et al., 2021).
@@
In their seminal paper, Vaswani et al. (2017) introduced the Transformer model, with the Self-Attention mechanism being its pivotal component. The basic idea behind the Self-attention module is illustrated in Figure A.1 and can be distilled as follows: the original input is fed into three different dense layers to generate three different datasets Query, Key and, Value. The Query and Key sets are used to calculate which part of the original input should be paid more attention to, then these attention weights are applied to the Value set to extract the salient features that deserve the most attention. The Self-attention module can be written as:
$\text{Self-attention(Q,K,V)}=\text{softmax}(\frac{QK^{T}}{\sqrt{d}})V$
@@
Prior to the introduction of the Transformer model, earlier NLP models that utilized the Self-attention mechanism still used RNNs for input representations (Cheng et al., 2016; Paulus et al., 2017). The Transformer model stands as the first model solely based on the Self-attention modules, forgoing the need for convolutional or recurrent layers. Since its inception, the Transformer model swiftly becomes the premier choice for NLP tasks, and myriad NLP models are built on the Transformer structure. Though the Transformer model was originally proposed to compete with other seq2seq models, it has been proven to be versatile and can be used in a wide range of modern deep learning applications. For example, the word embedding model employed in this paper is built on the encoder component of the Transformer model.
One of the most influential Transformer-based models is the Bidirectional Encoder Representations from Transformers (BERT) model built by Devlin et al. (2018). [^20] Another influential NLP model built on Transformer is Generative Pre-training Transformer (GPT), which is the backbone of Chat-GPT application. BERT is a pre-trained large language model, and it is trained on Wikipedia and BooksCorpus, collectively encompassing approximately 3.3 billion words. Not surprised, BERT obtains state-of-the-art results on multiple mainstream NLP tasks.
@@
In this study, we take the first approach and use the BERT model directly for word embedding, and the vectorized text content will be the input for a customized deep learning model. Word embedding represent words as vectors of real numbers but still capture these words’ meaning and semantic relationships. However, there is one caveat on applying word embedding on a long transcript document: most off-the-shelf pre-trained BERT models can only process up to 512 tokens at a time. This limitation stems from the quadratic complexity ($O(n^{2})$) of the self-attention layer in BERT concerning sequence length ($n$). Imposing this token cap prevents excessive memory consumption and computational overheads. There are several strategies that are commonly used to circumvent the limitation: (1) selectively processing the most informative portions of the text to alleviate computational burdens. For example, Chen et al. (2022) choose to process the beginning and end of news articles to extract the most informative texts. (2) dividing the original text into segments and processing them individually, as pursued by Wang et al. (2019).(3) introducing modified transformer models to accommodate longer texts (Rae et al., 2019; Kitaev et al., 2020; Ding et al., 2020).
@@
Appendix References
  • Chen et al. (2022) Chen, X., P. Cong, and S. Lv (2022). A long-text classification method of chinese news based on bert and cnn. IEEE Access 10, 34046–34057.
  • Cheng et al. (2016) Cheng, J., L. Dong, and M. Lapata (2016). Long short-term memory-networks for machine reading. arXiv preprint arXiv:1601.06733.
  • Devlin et al. (2018) Devlin, J., M.-W. Chang, K. Lee, and K. Toutanova (2018). Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805.
  • Ding et al. (2020) Ding, M., C. Zhou, H. Yang, and J. Tang (2020). Cogltx: Applying bert to long texts. Advances in Neural Information Processing Systems 33, 12792–12804.
  • Gentzkow et al. (2019) Gentzkow, M., B. Kelly, and M. Taddy (2019). Text as data. Journal of Economic Literature 57(3), 535–74.
  • Kitaev et al. (2020) Kitaev, N., Ł. Kaiser, and A. Levskaya (2020). Reformer: The efficient transformer. arXiv preprint arXiv:2001.04451.
  • Minaee et al. (2021) Minaee, S., N. Kalchbrenner, E. Cambria, N. Nikzad, M. Chenaghlu, and J. Gao (2021). Deep learning–based text classification: a comprehensive review. ACM computing surveys (CSUR) 54(3), 1–40.
  • Paulus et al. (2017) Paulus, R., C. Xiong, and R. Socher (2017). A deep reinforced model for abstractive summarization. arXiv preprint arXiv:1705.04304.
  • Rae et al. (2019) Rae, J. W., A. Potapenko, S. M. Jayakumar, and T. P. Lillicrap (2019). Compressive transformers for long-range sequence modelling. arXiv preprint arXiv:1911.05507.
  • Vaswani et al. (2017) Vaswani, A., N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, Ł. Kaiser, and I. Polosukhin (2017). Attention is all you need. Advances in neural information processing systems 30.
  • Wang et al. (2019) Wang, Z., P. Ng, X. Ma, R. Nallapati, and B. Xiang (2019). Multi-passage bert: A globally normalized bert model for open-domain question answering. arXiv preprint arXiv:1908.08167.
@@ Appendix B Results for Robustness Checks
Table B.1: Voting Results and Personal Experience
$v_{ij}$ (Probit Model)
(1) (2) (3)
Macro Factors
$T_{unemp}$ 0.0275-0.0275 $0.0049$ $0.0085$
$(0.1724)$ $(0.2984)$ $(0.1854)$
$D_{unemp}$ $0.1610$ 0.1444-0.1444 $0.2228$
$(0.1446)$ $(0.3071)$ $(0.1750)$
$T_{CPI}$ 0.1245-0.1245 $0.1946$ 0.2324-0.2324
$(0.1663)$ $(0.1690)$ $(0.2216)$
$D_{CPI}$ 0.1258-0.1258 $0.0386$ 0.0400-0.0400
$(0.2747)$ $(0.2647)$ $(0.4146)$
Interaction with Personal Experience
$T_{unemp}\times\text{Great Depression}$ $0.1370$
$(0.3846)$
$D_{unemp}\times\text{Great Depression}$ 0.0031-0.0031
$(0.4289)$
$T_{CPI}\times\text{Great Depression}$ $0.4767^{*}$
$(0.2551)$
$D_{CPI}\times\text{Great Depression}$ $0.3002$
$(0.4030)$
$T_{unemp}\times\text{Great Inflation}$ 0.0127-0.0127
$(0.3484)$
$D_{unemp}\times\text{Great Inflation}$ $0.4185$
$(0.3450)$
$T_{CPI}\times\text{Great Inflation}$ 0.2437-0.2437
$(0.2507)$
$D_{CPI}\times\text{Great Inflation}$ $0.0881$
$(0.4030)$
$T_{unemp}\times\text{WWII}$ 0.0539-0.0539
$(0.3356)$
$D_{unemp}\times\text{WWII}$ 0.3148-0.3148
$(0.3371)$
$T_{CPI}\times\text{WWII}$ $0.4902^{*}$
$(0.2735)$
$D_{CPI}\times\text{WWII}$ $0.1226$
$(0.4755)$
Fixed Effects Yes Yes Yes
Log Likelihood -328.5406 -328.7668 -327.3776
NN 2,539 2,539 2,539
  • Note: This table examines how members’ personal experience may interact with macroecomic situations and affect observed voting results. The sample period spans from the February 1986 FOMC meeting, when Tealbook core CPI data first became available, to the December 2018 meeting. $^{***}p<0.01$; $^{**}p<0.05$; $^{*}p<0.1$
@@ C.1 Hidden dissent inside FOMC members’ speeches
Utilizing the score $hd_{ij}^{speech}$ from members’ speeches, we can also construct the variable $HD_{i}^{speech}$ to measure the hidden dissent level prior to each FOMC meeting. These speeches are linked to meeting $i$ if delivered between meetings i1i-1 and $i$, and $HD_{i}^{speech}$ is calculated using the same method as $HD_{i}$. Summary statistics for these hidden dissent measures are reported in Table C.1.
@@ C.2 Do public speeches reveal hidden dissent in advance?
As discussed in subsection C.1, we feed FOMC members’ speeches into the deep learning model and generate two hidden dissent scores for each speech, $hd_{ij}^{speech,in}$ and $hd_{ij}^{speech,pre}$, depending on how to chose the reference point for chair. The first measure, $hd_{ij}^{speech,in}$, uses the chair’s words from the following meeting as the reference point. Since transcripts are subject to a 5-year delay and are not available in real time, the second measure, $hd_{ij}^{speech,pre}$, uses the chair’s speeches prior to the meeting, which are available in real time, as the reference point instead. For both measures, we compute $HD_{i}^{speech}$ for each meeting $i$ by averaging the hidden dissent scores for speeches delivered between meetings i1i-1 and $i$.